# Teaching Human Behavior Improves Content Understanding Abilities Of LLMs

### v1

@@ Abstract
Communication is defined as “Who says what to whom with what effect.” A message from a communicator generates downstream receiver effects, also known as behavior. Receiver behavior, being a downstream effect of the message, carries rich signals about it. Even after carrying signals about the message, the behavior data is often ignored while training large language models. We show that training LLMs on receiver behavior can actually help improve their content-understanding abilities. Specifically, we show that training LLMs to predict the receiver behavior of likes and comments improves the LLM’s performance on a wide variety of downstream content understanding tasks. We show this performance increase over 40 video and image understanding tasks over 23 benchmark datasets across both 0-shot and fine-tuning settings, outperforming many supervised baselines. Moreover, since receiver behavior, such as likes and comments, is collected by default on the internet and does not need any human annotations to be useful, the performance improvement we get after training on this data is essentially free-lunch. We release the receiver behavior cleaned comments and likes of 750k images and videos collected from multiple platforms along with our instruction-tuning data.
Equal Contribution. Contact ykumar@adobe.com for questions and suggestions.
Figure 1: The diagram depicts the five factors of communication in the context of an example YouTube video https://www.youtube.com/watch?v=p8ZYgb8Osek and where lies the free lunch. The receiver effect is not used while training Large Vision and Language Models. However, it contains many important signals that can help in understanding the content. The figure shows several comments containing temporal, cognitive, character, context, and user opinion information useful for understanding the video.
@@ 1 Introduction
Communication is defined by five factors: sender, message, channel, receiver, and behavior [71, 45, 46]. Lasswell encoded these five factors in the phrase, “Who says what to whom with what effect.” [71, 45, 46]. Human behavior occurs as a downstream artifact in the process of communication. Behavior is produced by the receiver as a response to the message sent by the sender. Being a downstream effect, behavior can help us infer important signals about the message itself. These signals, if properly harnessed, should be able to increase performance on the message understanding tasks popular in NLP and CV, like question answering, sentiment analysis, topic classification, etc. Despite this, behavior data is considered noise and is ignored while training large language models [15, 63] and also large vision and language models [54, 91]. In this paper, we explore this line of thought more.
@@
Humans produce two kinds of behavioral signals upon observing a message [11, 66]: perceptual signals and actions as behavior. Perceptual signals, like seeing, touching, and hearing, help a receiver primarily sense the world around her, ultimately guiding her actions. Actions are how a receiver acts on the outside world. The signals produced by the human receiver upon receiving a message carry information about the message itself (Fig. 1). For instance, the way a person’s eyes move while reading a book or how his heartbeat rises and falls upon watching a movie can help us extract useful information about the book or the movie. In this case, a rising heartbeat is often indicative of an exciting scene in the movie, and regressing while reading is indicative of important or confusing phrases [14]. Similarly, the actions a person performs after watching a movie, such as comments and likes, carry signals about the movie. Expanding on these ideas, prior literature has shown that harnessing perceptual signals, like eye movements, saliency, keystrokes, mouse movements, and FMRI, by modeling them together with content understanding tasks can improve both NLP and CV tasks. For instance, performance improvement is observed in tasks like visual and natural language question answering [62, 41, 74], text and image sentiment analysis [41, 10, 25], natural language inference [41], part-of-speech identification [8, 9], named entity recognition [31, 30], syntactic parsing [65], image captioning [21], visual object detection [77, 43].
@@
While the initial studies show that perceptual signals have much promise for improving downstream content understanding, they have a few significant issues due to which integrating human perception has not seen wide adoption in training LLMs. The human processing signals can only be collected in lab settings requiring specialized lab equipment and are thus expensive to collect and are limited in number. For example, the largest datasets containing the human processing signals are SALICON [36] and Cheng et al. [19] for visual saliency (10k images each), CELER [12] containing 28k sentences and human eye movements of 365 participants reading them, Dundee corpus [38] containing eye movements over 20 news articles, and Dhakal et al. [22] containing keystroke patterns over 1.5k sentences. Clearly, these datasets, while making important contributions, do not scale to the level at which today’s large language models are trained (trillions of natural language and image tokens). Further, being restricted to lab settings also limits the types of demographics reachable by the lab, thus limiting their representativeness of human diversity. Lab settings also lead to people changing their behavior in response to their awareness of being observed, an effect commonly called the Hawthorne effect [59, 70, 58]. On the other hand, actions (the other type of behavioral signals produced by a human receiver) are collected at a large scale in the form of digital analytics. Examples of this kind of data are likes, views, shares, comments, and purchase histories on images, tweets, videos, webpages, and other kinds of media. Action data has a much broader representation than is possible in lab settings and is much cheaper to collect than using specialized lab equipment. At the same time, actions have not been much investigated in the literature for their potential to improve downstream content understanding. Further, they are much more noisier than perceptual signals collected in the lab, owing to their collection in real-world settings.
@@
Therefore, in this paper, we make initial efforts to collect and understand digital analytics at scale with the aim of integrating them with LLMs to improve their downstream content understanding capabilities. We introduce methods for filtering and cleaning behavioral data and then propose tasks for large language and vision models, leading to improvements in language and visual content understanding tasks. For this, we look to Reddit and YouTube as two major sources of visual content and human behavior in the form of viewer comments, likes, and upvotes. From Reddit, we collect 5 million images and videos along with their upvotes and top-rated comments from two major subreddits (r/pics and r/videos). Similarly, from YouTube, we collect 2.2 million videos from 30000 channels along with their likes, views, and top user comments. After extensive filtering and cleaning, we are left with 730k samples of videos and images across the two platforms which we use for the next steps.
Figure 3: Dense caption generated by Behavior-LLaVA for the video of a Volkswagen ad. The original video is posted at URL: https://www.youtube.com/watch?v=kyuGXPNr-T0. The red-colored text highlights the most important aspects of the video captured by Behavior-LLaVA, demonstrating an understanding of aesthetics, characters, world knowledge, emotion, and spatial relationships. More such examples are given in Figs. 4, 5, 7, and Figs. 6, 8 for images and videos respectively.
After collecting user behavior over image and video content, next, we design tasks to teach large vision and language models (VLMs) to simulate user behavior. For this, we use an instruction fine-tuning format. Given a video or an image and the other metadata like time of post and channel, we ask the model to simulate user behavior of likes and comments. See Fig 2, Listing 1 for examples. We choose LLaMA-Vid [52] as our base model to teach it the user behavior. We call the resultant model Behavior-LLaVA (Large Language and Vision Assistant) [54]. Note that all the vision language models like LLaVA [54], MiniGPT-4 [91], Kosmos [32], BLIP [49], LLaMA-Vid [52] are trained on image-text instructions where text is some kind of annotations of the image, e.g., objects, object locations, captions, question-answers, etc. These instructions, since being sourced from either existing benchmark datasets such as MSCOCO [53] and Conceptual Captions [72] or being generated by stronger LLMs like GPT-4, require a significant amount of investment in terms of money for human annotation or generation through API. While building on top of their work, it is noteworthy that our model does not add any extra annotation or API cost to generate data. Rather, we rely on sampling data from digital analytics databases, which are anyway being collected for various reasons, and use them to improve VLM capabilities. Therefore, any potential improvements brought by integrating action data with VLMs do not require a significant capital investment.
We start with the trained LLaMA-Vid model. The LLaMA-Vid model uses two tokens to represent each video, which they call content and context tokens. While the context token encodes the overall image context based on user input, the content token encapsulates visual cues in each frame. For learning context tokens, the model uses attention queries that interact with previously generated image features in the designed attention module. To generate content tokens, the image features are average pooled. This dual-token strategy significantly reduces the number of tokens needed to represent videos, thus enabling the model to scale to longer (hour-long) videos. To better support hour-long videos, LLaMA-Vid was trained on a 9k movie-level conversation instruction set containing plot reasoning and detail understanding questions. Taking the base LLaMA-Vid model, we finetune it further on behavioral data. We teach the model a single instruction format given in Listing 1, where we teach it that given a media file (image or a video), it has to simulate the likes and comments on the media. We do this on a 730k instructions set created by us using Reddit and YouTube videos and images (§2). Next, we test the resultant model which we call Behavior-LLaVA, on a hierarchy of tasks in both zero-shot and finetuned settings. We include tasks based on low-level content understanding, like object and activity detection, and also on high-level understanding, like topic, sentiment, emotion, and persuasion classification. We also test the model on even higher-level tasks like memorability simulation and YouTube dislike/like simulation. We compare Behavior-LLaVA against the base model LLaMA-Vid and other state-of-the-art approaches reported on each test task (Tables 1-8).
@@
1) We explore the idea of learning human behavior, resulting in better content understanding. We test this for action-level behavior data corresponding to receiver comments and likes. We collect a dataset consisting of 400k images and 330k videos along with their receiver comments and likes. Then, LLaMA-Vid is trained for the task of predicting receiver comments and likes given a media (a video or an image) (Listing 1). We show that using this simple task formulation and behavioral data collected in the wild, results in performance improvement over a hierarchy of tasks. We get improvements over the base LLaMA-Vid across 40 tasks over 23 benchmark datasets in both zero-shot and fine-tuned settings. We show this over low-level content understanding tasks like object and activity recognition and also over high-level tasks like topic and emotion detection. Further, interestingly, none of the tasks tested show a decrease in performance by training LLaMA-Vid on this additional task. Through this, we propose a scalable approach to increase the content understanding abilities of VLMs, requiring minimal cost and no architectural changes.
2) To disentangle the effect of training LLaVA-Vid on the additional 730k video and image samples from the effect of training on behavior data over those samples, we train LLaVA-Vid on 730k samples using the original task of LLaVA-Vid. We call this model Ad-LLaVA. We show that Ad-LLaVA shows slight performance improvements over LLaVA-Vid; however, Behavior-LLaVA still performs better than both Ad-LLaVA and LLaMA-Vid, thus highlighting the importance of behavior data and instruction fine-tuning on behavior data.
3) We also show an ablation of Behavior-LLaVA across different kinds of behavior. We try out the perception behavior of saliency prediction over images and two types of action-level behavior (comments and likes) over images and videos. We find that perception-level behavior does not result in significant performance improvements; however, action-level behavior shows improvements across all the tasks. We posit that one reason for this could be due to the scale for which action-level data is available (Table 9). While perception behavior is mostly collected in lab settings, action-level behavior data can be collected in a scalable manner automatically and cheaply.
@@ 2.1 BLIFT Dataset
Given the abundance of media and behavioral data and its accessibility, our data collection relies on two primary sources: Reddit and YouTube. These platforms share similarities in terms of hosting media content (images and videos) and user engagement metrics such as Reddit upvotes and comments, and YouTube likes, views, and comments. Here, we delineate the process involved in constructing the instruction fine-tuning dataset, which we term as the Behavior-LLaVA Instruction Fine-Tuning (BLIFT) dataset.
@@ 2.1.1 Data from Reddit
To collect a substantial corpus of diverse images and videos, we targeted two specific subreddits, namely r/pics and r/videos. Established over 15 years ago, these subreddits had a user base exceeding 20 million during the data collection period, with an average of over 5,000 online users concurrently. Notably, due to stringent content moderation guidelines [69, 2, 3, 4] and the exclusive focus of these subreddits on media content, they offer a rich variety of content devoid of thematic biases. Our data collection spans until January 2022, during which the activity on these subreddits witnessed a notable decline following several policy adjustments and user protests [29, 24].
To ensure data quality and relevance, we executed a series of filtering steps on the posts and comments from these subreddits. Initially, we excluded posts predating February 2015 from r/pics, coinciding with the implementation of a rule requiring images without digital/overlay text [2, 3]. This filtering step resulted in the exclusion of 3.1 million images and 2 million videos. Subsequently, considering the sustained popularity of both subreddits, with rankings within the top 20 since 2017 and consistent membership exceeding 20 million, we confined our dataset to posts from January 2018 onwards. This selection process yielded 1.4 million images and 1.1 million videos.
@@
<SYSTEM>You are an AI visual assistant, you are given a detailed description of a media, followed by the actual media. Answer all questions as you are seeing the media<SYSTEM><USER>The video advertisement is titled Gatorade | Make Your Own Footsteps with Suni Lee for the brand Gatorade. The audio in the ad says ”[ASR HERE …]”.
Analyze this video deeply, then write scene by scene description of the video and answer the following questions. What percentage of viewers would like this video, and what would be the top-5 popular comments on this video?.
<video>.. </video>
<USER><ASSISTANT>
The scene by scene descriptions are:
Scene 1: The scene shows a woman looking off into the distance with an orange line going around the outside of the screen. The foreground colors of the scene are Black,Mud_Green,Gray,Dark_Gray, and the background colors are Dark_Brown,Black,Dark_Gray. The dominant tone of the scene is neutral. This scene is categorized by the tags cosmetic, eyebrow, face, girl, ponytail, stand, string, woman.
Scene 2: The scene shows a woman balancing on a skate board in a yard. The foreground colors of the scene are Black,Mud_Green,Dark_Gray,Olive, and the background colors are Black,Dark_Gray,Gray,Dark_Brown. The dominant tone of the scene is neutral. This scene is categorized by the tags athletic, balance, beam, car, girl, house exterior, hurdle, jog, legging, plank, rail, seesaw, woman, yard.
Scene 3: The scene shows a girl jumping over a wooden ramp in the backyard. The foreground colors of the scene are Black,Dark_Gray,Gray,Dark_Blue, and the background colors are Dark_Brown,Dark_Blue,Purple,Dark_Pink,Brown. The dominant tone of the scene is neutral. This scene is categorized by the tags arm, backyard, balance, diving board, girl, house exterior, ledge, plank, purple, rail, shirt, short, stand, stretch, woman, yard.
@@
1. Wow. Love it. Shes such an inspiration to the the next generation as well as everyone.
2. Inspiring and great story behind this commercial. Builds meaning and purpose in the hearts and mind of youth. Its been awhile since good meaningful ads have been made.
3. Shes an inspiration to the world thanks to her my niece learning Gymnastics hopefully someday, she is an inspiration to other youngers as how suni is an inspiration to everyone
4. Chills watching this. Such an inspiration
5. Yooooo, this is straight up !
<ASSISTANT>

### v3

# Veeky Baths [Uncaptioned image], Rajiv Ratn Shah [Uncaptioned image], Changyou Chen [Uncaptioned image], Balaji Krishnamurthy [Uncaptioned image]

@@ Abstract
Communication is defined as “Who says what to whom with what effect.” A message from a communicator generates downstream receiver effects, also known as behavior. Receiver behavior, being a downstream effect of the message, carries rich signals about it. Even after carrying signals about the message, the behavior signal is often ignored while training vision language models. We show that training VLMs on receiver behavior can actually help improve their content-understanding abilities. We demonstrate that training VLMs to predict receiver behaviors, such as likes, comments, and replay graphs, which are available at scale, enhances the VLM’s performance across a broad range of downstream content understanding tasks. We show this performance increase over 6 types of behavior, 46 different tasks covering image, video, text and audio over 26 benchmark datasets across both 0-shot and fine-tuning settings, outperforming many supervised baselines on diverse tasks ranging from emotion recognition to captioning by upto 150%. We note that since receiver behavior, such as likes, comments, and replay graphs, is collected by default on the internet and does not need any human annotations to be useful, the performance improvement we get after training on this data is essentially free-lunch. We also release BLIFT, our Behaviour-LLaVA IFT dataset comprising of 730k images and videos with their receiver behavior collected from multiple platforms on which we train our models to achieve this.
Figure 1: The diagram depicts the five factors of communication in the context of an example YouTube video https://www.youtube.com/watch?v=eT8hO4e2iTM and where lies the free lunch. The receiver effect is not used while training Large Vision and Language Models. However, it contains many important signals that can help in understanding the content. The figure shows several comments containing temporal, cognitive, character, context, and user opinion information useful for understanding the video.
{NoHyper}Equal Contribution. Contact behavior-in-the-wild@googlegroups.com for questions and suggestions.
@@ 1 Introduction
Communication is defined by five factors: sender, message, channel, receiver, and behavior (Shannon & Weaver, 1949; Lasswell, 1948; 1971). Lasswell (1948) encoded these five factors in the phrase, “Who says what to whom with what effect.” Human behavior occurs as a downstream artifact in the process of communication. Behavior is produced by the receiver as a response to the message sent by the sender. Being a downstream effect, behavior can help us infer important signals about the message itself. These signals, if properly harnessed, should be able to increase performance on the message understanding tasks popular in NLP and CV, like question answering, sentiment analysis, topic classification, etc. Despite this, behavior data is considered noise and is ignored while training large language models (Biderman et al., 2022; Penedo et al., 2023) and also large vision and language models (Liu et al., 2023a; Zhu et al., 2023). In this paper, we explore this line of thought more.
@@
Humans produce two kinds of behavioral signals upon observing a message (Bertenthal, 1996; Prinz, 1997): perceptual signals and actions as behavior. Perceptual signals, like seeing, touching, and hearing, help a receiver primarily sense the world around her, ultimately guiding her actions. Actions are how a receiver acts on the outside world. The signals produced by the human receiver upon receiving a message carry information about the message itself (Fig. 1). For instance, if a person’s heartbeat rises upon watching a movie scene, it can help us infer that perhaps the scene was an exciting scene (Dzedzickis et al., 2020). Similarly, regressing while reading is indicative of important or confusing phrases (Bicknell & Levy, 2011). In these cases, perception behavior helps us derive inferences about content. In a similar vein, the actions a person performs after watching a movie, such as comments and likes, carry signals about the movie (Fig. 1, 2).
Expanding on these ideas, prior literature has shown that harnessing perceptual signals, like eye movements, saliency, keystrokes, mouse movements, and FMRI, by modeling them together with content understanding tasks can improve both NLP and CV tasks. For instance, integration with perception signals causes performance improvement in tasks like visual and natural language question answering (Patro & Namboodiri, 2018; Khurana et al., 2023; Sood et al., 2020), text and image sentiment analysis (Khurana et al., 2023; Barrett et al., 2018; Fan et al., 2018), natural language inference (Khurana et al., 2023), part-of-speech identification (Barrett et al., 2016a; b), named entity recognition (Hollenstein et al., 2019; Hollenstein & Zhang, 2019), syntactic parsing (Plank, 2016), image captioning (Cornia et al., 2018), and visual object detection (Wang et al., 2018; Kruthiventi et al., 2016).
@@
While the initial studies show that perceptual signals have much promise for improving downstream content understanding, they have a few significant issues due to which integrating human perception has not seen wide adoption in training LLMs. These perceptual signals can only be collected in lab settings requiring specialized lab equipment and are thus expensive to collect and thus are also limited in number. For example, the largest datasets containing the human processing signals are SALICON (Jiang et al., 2015) and Cheng et al. (2014) for visual saliency (10k images each), CELER (Berzak et al., 2022) and Dundee corpus (Kennedy et al., 2013) containing eye movements over 28k sentences and 20 news articles respectively, and Dhakal et al. (2018) containing keystroke patterns over 1.5k sentences. Clearly, these datasets, while making important contributions, do not scale to the level at which today’s large language models are trained (trillions of natural language and image tokens).
On the other hand, actions (the other type of behavioral signals produced by a human receiver) are collected at a large scale in the form of digital analytics. Examples of this kind of data are likes, views, shares, comments, and purchase histories on images, tweets, videos, webpages, and other kinds of media. Action data has a much broader representation than is possible in lab settings, is available on more diverse content, and is much cheaper to collect than using specialized lab equipment. At the same time, actions have not been much investigated in the literature for their potential to improve downstream content understanding.
@@
Therefore, in this paper, we make initial efforts to collect and understand digital analytics at scale with the aim of integrating them with VLMs to improve their downstream content understanding capabilities. We introduce methods for filtering and cleaning behavioral data and then propose tasks for large language and vision models, leading to improvements in language and visual content understanding tasks. For this, we look to Reddit and YouTube as two major sources of visual content and human behavior in the form of viewer comments, likes, replay graphs, and upvotes. From Reddit, we collect 5 million images and videos along with their upvotes and top-upvoted comments from two major subreddits (r/pics and r/videos). Similarly, from YouTube, we collect 2.2 million videos from 30000 channels along with their likes, views, replay graphs, and top user comments. After extensive filtering and cleaning, we are left with 730k samples of videos and images across the two platforms which we use for the next steps.
After collecting user behavior over image and video content, we design tasks to teach large vision and language models (VLMs) to simulate user behavior. For this, we use an instruction fine-tuning format. Given a video or an image and the other metadata like time of post and channel, we ask the model to simulate user behavior of likes and comments. See Fig 2, Listing 1 for examples. We choose LLaMA-Vid (Li et al., 2023b) as our base model to teach it the user behavior. We call the resultant model Behavior-LLaVA (Large Language and Vision Assistant) (Liu et al., 2023a). We test Behavior-LLaVA on a diverse variety of tasks, evaluating its capabilities on image, video, text, and audio understanding tasks. We compare Behavior-LLaVA against its base model, LLaMA-Vid, and other supervised baselines. Further, to show the impact of behavior, we train another version of LLaMA-Vid, where we train it on the same set of videos and images as Behavior-LLaVA but do not include behavior information. We call this model Ad-LLaVA.
Figure 3: Behaviour-LLava achieves much higher zero-shot performance compared to Ad-LLaVA and the base model LLaMA-VID across a diverse suite of image, video, and audio benchmarks.
@@
1) Behavior-LLaVA Instruction Fine-Tuning: We explore the idea of learning human behavior, resulting in better content understanding. We test this for action-level behavior data such as receiver comments, likes, and replay graphs. We collect a dataset called BLIFT, consisting of 400k images and 330k videos, along with their receiver behavior. Then, LLaMA-Vid is trained for the task of predicting receiver comments and upvotes given a media (a video or an image) (Listing 1). We show that using this simple task formulation over behavioral data collected in the wild, results in performance improvement over a hierarchy of tasks. We get improvements over the base LLaMA-Vid across 46 tasks over 26 benchmark datasets in both zero-shot and fine-tuned settings. We show this over low-level content understanding tasks like object and activity recognition and also over high-level tasks like topic and emotion detection. Through this, we propose a scalable approach to increase the content understanding abilities of VLMs, requiring minimal cost and no architectural changes.
2) AdLLaVA: Disentangling the effect of content and behaviour: To disentangle the effect of training LLaMA-Vid on additional image and video data from the effect of training on behavior data, we train LLaMA-Vid on BLIFT’s videos and images without including behavior. We call this model Ad-LLaVA. We show that Ad-LLaVA shows equivalent performance as its base model LLaVA-Vid; however, Behavior-LLaVA performs better than both Ad-LLaVA and LLaMA-Vid, thus highlighting the importance of behavior data and instruction fine-tuning on behavior data.
3) Perception vs Action:We also show an ablation of Behavior-LLaVA across different kinds of behavior. We try out the perception behavior of saliency prediction over images and five types of action-level behavior over images and videos. We find that perception-level behavior does not result in significant performance improvements; however, action-level behavior shows improvements across all the tasks. We posit that one reason for this could be due to the scale for which action-level data is available (Table 11). While perception behavior is mostly collected in lab settings, action-level behavior data is diverse, can be collected in a scalable manner automatically and cheaply.
@@ 2.1 BLIFT Dataset
Given the abundance of media and behavioral data and its accessibility, our data collection relies on two primary sources: Reddit and YouTube. These platforms share similarities in terms of hosting media content (images and videos) and providing user engagement metrics in the form of Reddit upvotes and comments, and YouTube likes, views, comments, and replay graphs. Here, we delineate the process involved in constructing the instruction fine-tuning dataset, which we term as the Behavior-LLaVA Instruction Fine-Tuning (BLIFT) dataset.
@@ 2.1.1 Data from Reddit
To collect a substantial corpus of diverse images and videos, we targeted at two specific subreddits, namely r/pics and r/videos. Established over 15 years ago, these subreddits had a user base exceeding 20 million during the data collection period, with an average of over 5,000 online users concurrently. Notably, due to stringent content moderation guidelines (Reddit, Inc., 2024; red, 2024a; b; c) and the exclusive focus of these subreddits on media content, they offer a rich variety of content devoid of thematic biases. Our data collection spans until January 2022, during which the activity on these subreddits witnessed a notable decline following several policy adjustments and user protests (Hern, 2023; Economist, 2023).
To ensure data quality and relevance, we executed a series of filtering steps on the posts and comments from these subreddits. Initially, we excluded posts predating February 2015 from r/pics, coinciding with the implementation of a rule requiring images without digital/overlay text (red, 2024a; b). This filtering step resulted in the exclusion of 3.1 million images and 2 million videos. Subsequently, considering the sustained popularity of both subreddits, with rankings within the top 20 since 2017 and consistent membership exceeding 20 million, we confined our dataset to posts from January 2018 onwards. This selection process yielded 1.4 million images and 1.1 million videos.
@@
<SYSTEM>You are an AI visual assistant. You are given a detailed description of a media, followed by the actual media. Answer all questions as if you are seeing the media.</SYSTEM>
<USER>
The video advertisement is titled "Gatorade | Make Your Own Footsteps with Suni Lee" for the brand Gatorade. The audio in the ad says "[ASR HERE ...]". Analyze this video deeply, then write scene by scene description of the video and answer the following questions. What percentage of viewers would like this video, and what would be the top-5 popular comments on this video? What would the replay graph values for each scene be?
<video>...</video>
</USER>
<ASSISTANT>
The scene-by-scene descriptions are:
Scene 1: The scene shows a woman looking off into the distance with an orange line going around the outside of the screen. The foreground colors of the scene are black, mud green, gray, dark gray, and the background colors are dark brown, black, dark gray. The dominant tone of the scene is neutral. This scene is categorized by the tags: cosmetic, eyebrow, face, girl, ponytail, stand, string, woman.
Scene 2: The scene shows a woman balancing on a skateboard in a yard. The foreground colors of the scene are black, mud green, dark gray, olive, and the background colors are black, dark gray, gray, dark brown.The dominant tone of the scene is neutral. This scene is categorized by the tags: athletic, balance, beam, car, girl, house exterior, hurdle, jog, legging, plank, rail, seesaw, woman, yard.
Scene 3: The scene shows a girl jumping over a wooden ramp in the backyard. The foreground colors of the scene are black, dark gray, gray, dark blue, and the background colors are dark brown, dark blue, purple, dark pink, brown.
$>$ BEHAVIOR $<$
@@
1. "Wow. Love it. Shes such an inspiration to the next generation as well as everyone."
2. "Inspiring and great story behind this commercial. Builds meaning and purpose in the hearts and minds of youth. Its been a while since good, meaningful ads have been made."
3. "Shes an inspiration to the world. Thanks to her,my niece is learning gymnastics. Hopefully someday, she is an inspiration to others as Suni is an to everyone"
4. "Chills watching this. Such an inspiration."
5. "Yooooo, this is straight up!"
The replay values for each scene would be:
Scene 1: 0.06
Scene 2: 0.23
Scene 3: 0.38
</ASSISTANT>
Figure 4: Behavior Instruction fine-tuning template for the video: https://www.youtube.com/watch?v=BKPQkjRF4yY